OpenStudy (anonymous):

Check my answer please I will give a medal Solve the equation. (equation below)

OpenStudy (anonymous):

\[\frac{ d }{ 3 } + 10 = 7\]

OpenStudy (anonymous):

My answer: d = -9

OpenStudy (anonymous):

@thomaster @mathstudent55

OpenStudy (anonymous):

d/3+10=7; d/3=-3;d=-9

OpenStudy (anonymous):

Thanks